Lesson 14 Stoichiometry Problems

  1. The compound KClO3 decomposes according to the following equation: 2KClO3 = 2KCl + 3O2

a. What is the mole ratio of KClO3 to O2 in this reaction?

b. How many moles of O2 can be produced by letting 6.0 moles of KClO3 react based on the above equation?

c. How many molecules of oxygen gas, O2, are produced in question 1b?

2.Magnesium combines with chlorine, Cl2, to form magnesium chloride, MgCl2, during a synthesis reaction.

a.Write a balanced chemical equation for the reaction.

b. How many moles of magnesium chloride can be produced with 3 moles of chlorine

3.Ethanol burns according to the following equation. If 15.0 mol of C2H5OH burns this way, how many moles of 

oxygen are needed?

C2H5OH + 3O2 = 2CO2 + 3H2O


1
Expert's answer
2021-03-22T06:29:10-0400

Solution:

Question 1:

The balanced chemical equation:

2KClO3 → 2KCl + 3O2

According to the equation above: n(KClO3)/2 = n(KCl)/2 = n(O2)/3


(a): Moles of KClO3 : Moles of O2 = 2 : 3


(b): n(O2) = 3 × n(KClO3) / 2 = (3 × 6.0 mol) / 2 = 9.0 mol

Moles of O2 = 9.0 mol


(c): One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Number of O2 molecules = (9.0 mol O2) × (6.022×1023 molecules / 1 mol O2) = 5.42×1024 O2 molecules

Number of O2 molecules = 5.42×1024 molecules


Answers to Q1:

a) Moles of KClO3 : Moles of O2 = 2:3; b) Moles of O2 = 9.0 mol; c) 5.42×1024 molecules.


Question 2:

(a): The balanced chemical equation: Mg + Cl2 → MgCl2


(b): According to the equation above: n(Mg) = n(Cl2) = n(MgCl2)

Hence,

n(MgCl2) = n(Cl2) = 3.0 mol

Moles of MgCl2 = 3.0 mol


Answers to Q2: a) Mg + Cl2 → MgCl2; b) Moles of MgCl2 = 3.0 mol.


Question 3:

The balanced chemical equation:

C2H5OH + 3O2 → 2CO2 + 3H2O

According to the equation above: n(C2H5OH) = n(O2)/3 = n(CO2)/2 = n(H2O)/3

Hence,

n(O2) = 3 × n(C2H5OH) = 3 × 15.0 mol = 45.0 mol

Moles of O2 = 45.0 mol


Answer to Q3: Moles of O2 = 45.0 mol.